OpenStudy (whatever):

a 2 digit number can be written in the form 10x+y.if the digits are reserved, the new number is 10y+x.to solve coin problems, you can write one equation for the value of the coins. a collection of dimes and quarters iss worth $25. there are 40 more domes than quarters how many of each are there? x: dimes y: quarters 0.1x+0.25y=25 and y+40=x (ive tried doing this but i have feeling that i am doing it wrong so please someone help me and explain it to me and show steps?)

OpenStudy (anonymous):

\[0.1(y+40)+0.25y=25\] should be enough to help you through this one.

OpenStudy (whatever):

okay thanks

OpenStudy (whatever):

wait i dont get it?

OpenStudy (anonymous):

okay, so \[0.1y+4+0.25y=25\] \[0.35y+4=25\] \[0.35y=21\] \[y=\frac{21}{0.35}=60\]

OpenStudy (anonymous):

medals don't hurt, ye know...